# Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany Now at ]Max-Planck-Institut für Gravitationsphysik (Albert-Einstein-Institut) and Leibniz Universität Hannover, 30167 Hannover, Germany

### v1

@@ IV.3.3 Field Overlap
Finally, it should be noted that in both runs, $\eta_{xy,{\rm veto}}^{2}$ and $\eta_{{\rm veto}}^{2}$ agree very with the results of the spatial overlap and total overlap measured at the science detector and discussed in Section IV.2.4.
@@ Acknowledgments
The work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and - grant number WI 1643/2-1, the Partnership for Innovation, Education and Research (PIER) of DESY and Universität Hamburg under PIER Seed Project - PIF-2022-18, the German Volkswagen Stiftung, the National Science Foundation - grant numbers PHY-2110705 365 and PHY-1802006, the Heising Simons Foundation - grant numbers 2015-154 and 2020-1841, the UK Science and Technologies Facilities Council - grant number ST/T006331/1.
@@ A.2 Changes in Open Shutter Data
(a) Magnitude squared changes open shutter $\rm S_{{}_{\perp}}$
(b) Angular changes open shutter $\rm S_{{}_{\perp}}$
(c) Magnitude squared changes open shutter $\rm S_{{}_{\parallel}}$
(d) Angular changes open shutter $\rm S_{{}_{\parallel}}$

### v2

@@ IV.2.5 Field Overlap Results
Finally, it should be noted that in both runs, like $\eta_{xy,{\rm veto}}^{2}$, $\eta_{{\rm veto}}^{2}$ agree very with the results of the spatial overlap and total overlap measured at the science detector and discussed in Section IV.2.4.
IV.2.6 Optimizing the Spatial Overlap
@@ Acknowledgments
We would like to thank the ALPS II collaboration, especially David Reuther, Sandy Croatto, and Sven Karstensen, the FTX group, especially, Karsten Gadow and Maik Dinter, the MVS group, especially Antonio de Zubiaurre-Wagner, MKS, FEB, MPC, MPS, ZM1 and ZM5, and the MEA2 and MEA5 groups at DESY as well as the FH mechanical workshop, for their support during the commissioning and operation of the experiment. We are also grateful to Manuel Meyer for the many valuable discussions and his feedback regarding the manuscript.
We acknowledge the support of the National Science Foundation (Grant No. 1802006), of the Heising-Simons Foundation (Grant No. 2015-154 and 2020-1841), of the Deutsche Forschungsgemeinschaft through project grant WI 1643/2-1 and EXC 2121 “Quantum Universe” – 390833306, of the Science and Technology Facilities Council (UK), grants ST/T006331/1 and ST/Y004515/1, as well as support by the German Volkswagen Stiftung and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program Grant agreement No. 948689.
@@ A.2 Changes in Open Shutter Data
(a) Magnitude squared changes open shutter ($\rm S_{{}_{\perp}}$)
(b) Angular changes open shutter ($\rm S_{{}_{\perp}}$)
(c) Magnitude squared changes open shutter ($\rm S_{{}_{\parallel}}$)
(d) Angular changes open shutter ($\rm S_{{}_{\parallel}}$)